OpenStudy (anonymous):

An empty box is shaped like a rectangular prism. The box has a base area of 9/10 square foot and a height of 1/3 foot. How much packing material is required to fill the box? A. 3/10 ft^3 B. 1/10 ft^3 C. 1/3 ft^3 D. 2/3 ft^3

OpenStudy (anonymous):

To find the volume of the prism you have to multiply length(L) width(W) and height(H) You are already given the height so H=1/3 And the base area means they did L*W 9/10*1/3=3/10 the answer is A

OpenStudy (anonymous):

Do you understand it?

OpenStudy (anonymous):

yes thank you im just horrible with fractions